OpenStudy (anonymous):

does this equation open left or right? x=1/4y^2

OpenStudy (anonymous):

I am confused b/c the question is basically the equation of a horizontal parabola

OpenStudy (anonymous):

Yes, which way are the arrows pointing to is the question.

OpenStudy (anonymous):

so "p" = 1 making it greater than 0 so it must be going right?

OpenStudy (anonymous):

Well, just think about this: On a y=x^2 graph, it opens up. On a y=-x^2 graph, it opens down. You have x=y^2 here. Which side is going in the positive direction? The right. So you are correct but you don't have to play with numbers necessarily. This is easier as well.